A 100-m-long copper wire of radius 0.12 mm is connected across a 1.5-V battery has 9.4 x 1019 electrons passing a cross section of the wire every 3.0 s. Calculate the resistance of this wire.


1
Expert's answer
2022-05-31T16:11:38-0400
"I=\frac{q}{t}=\frac{9.4*10^{19}*1.6*10^{-19}}{3.0}=5.0\:\rm A"

"R=\frac{V}{I}=\frac{1.5}{5.0}=0.3\:\Omega"